### v1

# Towards Scalability and Extensibility of Query Reformulation Modeling in E-commerce Search

Abstract.
Customer behavioral data significantly impacts e-commerce search systems. However, in the case of less common queries, the associated behavioral data tends to be sparse and noisy, offering inadequate support to the search mechanism. To address this challenge, the concept of query reformulation has been introduced. It suggests that less common queries could utilize the behavior patterns of their popular counterparts with similar meanings. In Amazon product search, query reformulation has displayed its effectiveness in improving search relevance and bolstering overall revenue. Nonetheless, adapting this method for smaller or emerging businesses operating in regions with lower traffic and complex multilingual settings poses the challenge in terms of scalability and extensibility. This study focuses on overcoming this challenge by constructing a query reformulation solution capable of functioning effectively, even when faced with limited training data, in terms of quality and scale, along with relatively complex linguistic characteristics. In this paper we provide an overview of the solution implemented within Amazon product search infrastructure, which encompasses a range of elements, including refining the data mining process, redefining model training objectives, and reshaping training strategies. The effectiveness of the proposed solution is validated through online A/B testing on search ranking and Ads matching. Notably, employing the proposed solution in search ranking resulted in 0.14% and 0.29% increase in overall revenue in Japanese and Hindi cases, respectively, and a 0.08% incremental gain in the English case compared to the legacy implementation; while in search Ads matching led to a 0.36% increase in Ads revenue in the Japanese case.
E-commerce search, Query reformulation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 30th SIGKDD Conference on Knowledge Discovery and Data Mining - Applied Data Science Track; August 25–29, 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Computing methodologies Artificial intelligence
1. Introduction
Search ranking and matching in E-commerce heavily rely on historical customer behavior like clicks and purchases. These actions genuinely reflect customer preferences for products associated with specific queries, offering valuable insights for refining the search system (Agichtein et al., 2006). The refinements based on customer behavior further enables the formation of a beneficial feedback loop that intertwines increased customer engagement with elevated search experience. However, challenges arise in activating this loop, particularly for queries with less popular or newly emerging shopping intents, which are categorized as ”tail queries” (Goel et al., 2010).
An intuitive strategy to overcome this challenge involves enabling tail queries to benefit from the behavioral data of more popular queries that convey similar semantic meanings (Huang et al., 2023; Yin et al., 2016; Grbovic et al., 2015). This approach is implemented through query reformulation (QR), which creates linkages between queries grounded in their semantic similarity. Despite substantial efforts to apply sophisticated techniques in optimizing QR solutions, there remains a gap in achieving the robustness required for industrial QR practices.
In the realm of E-commerce, we consider the robustness of a QR solution from two key angles: scalability and extensibility. Scalability pertains to the size and quality of data required to support the effective model training. It is widely recognized that having a sufficient amount of high-quality training data can significantly ease the intricacies of the training process and reduce system complexity. However, for smaller markets or niche business fields, customer behavior signals often tend to be both sparse and noisy, making it challenging to construct a high-quality training dataset. Extensibility decides whether the QR approach can be applied across a broad spectrum of linguistic scenarios, including those with more intricate text patterns (take Japanese, for instance; unlike Indo-European languages, it comprises a mix of ideographic and phonetic characters, resulting in various script types, such as Kanji, Hiragana, Katakana, and Romaji, that may convey the same semantic meaning but are used in vastly different ways in diverse contexts). As such, our central focus in this paper is understanding how to build a QR solution capable of functioning effectively, even when faced with limited training data, in terms of quality and scale, along with relatively complex linguistic characteristics.
The primary challenge in achieving robustness of QR stems from its reliance on customer behavior data - the development of the QR framework is susceptible to behavioral noise caused by uneven query exposure to customers (Szpektor et al., 2011), evident in the form of false positive/negative query-query pairs within the training/evaluation data extracted from historical customer behavior. This noise are particularly prevalent with cold-start queries and those with limited association to related products due to system defects, which are exactly part of QR’s target. In light of this challenge, we have refined core concepts and components of the legacy reformulation modeling (Huang et al., 2023). Our objective is to extend the benefit of the QR framework to smaller, non-English markets in addition to expand the benefit in major English markets. This paper presents a comprehensive overview of these improvements. Specifically, we encapsulate the finding and contribution of this study as follows,
  • We introduce novel measurements for query-query relationship mining with behavioral data, specifically designed to reduce biases stemming from false positives during downstream model training.
  • We refine model training objectives by introducing a sample importance factor, and also integrate hard negative mining into the model training pipeline to improve recall and precision, respectively at the retrieval and re-ranking stage of QR.
  • Through comprehensive offline analyses, we illustrate the proposed approaches effectively mitigate behavioral noise, thereby enhancing offline recall@100 by 35.2% and NDCG@3 in the hard negative scenario by 94.4% compared with the legacy QR modeling. Furthermore, results from human audits indicate that these approaches markedly strengthen the model’s ability to discern between relevant and challenging irrelevant query pairs, resulting in a noteworthy improvement in classification AUC from 0.51 to 0.79.
  • We executed online A/B tests in two distinct use cases of QR, including search ranking and search Ads matching, within Amazon product search. The experiment led to a noteworthy overall search revenue gain of +0.14% and +0.29% in Japanese and Hindi cases, respectively, and a +0.08% incremental revenue gain in English case compared to the legacy implementation. The experiment also led to a +0.36% increase in Ads revenue in the Japanese case. We also observed that the legacy QR solution that effectively works in major markets under English context showed no revenue gains in smaller markets with non-English language. This comparison confirmed the existence of scalability and extensibility issues, and also conclusively demonstrated the effectiveness of the proposed technique in ensuring robustness of QR.
2. Related work
Web service developers have dedicated substantial effort to query reformulation in order to address the inefficiencies of information retrieval systems (Carpineto and Romano, 2012).
Query-query relationship mining The basis of query reformulation lies in comprehending the connections between distinct queries. Previous research has explored various approaches, such as utilizing Bayesian perspectives (Antonellis et al., 2008; Cui et al., 2002; Fonseca et al., 2005) and graph-based approaches (Boldi et al., 2008, 2009; Beeferman and Berger, 2000) to mine query-query correlations leveraging query-document click-through data, which lay the groundwork for query rewriting using customer interaction data. This type of methods has been further enhanced with entropy information to counteract the bias towards popular queries (Deng et al., 2009). In this work, we also use query-document data for query similarity mining, but adopt new metrics based on distributional divergence customized for different training stages in QR. Another thread of work (Ozertem et al., 2012; Cao et al., 2021) utilized search log data to establish query-to-query affinity through session-based concepts. Yet, this approach suffers from rule-based restrictions and has a limited usage.
Query reformulation pipeline With the advent of transformer-based text encoders, there has been a paradigm shift toward leveraging the semantic information encoded within queries for rewrite tasks (Chen et al., 2020; Cho et al., 2021), employing concepts like semantic textual similarity and dense text retrieval. This entails generating representative query embeddings, enabling the establishment of query-to-query relationships on a large scale in a computationally efficient manner. While there has been a notable effort to develop QR modeling within E-commerce applications (Li et al., 2022; Qiu et al., 2021), these endeavors did not specifically focus on enhancing the behavioral signal of less common queries, and thus, the design of these models was not tailored accordingly. A previous work (Huang et al., 2023) bridges this gap for Amazon product search with a query reformulation pipeline, comprising two components: 1) A Query-to-Query retrieval model, designed to identify a multitude of candidate queries that share a similar shopping intent with a given query; 2) A Query-to-Query re-ranking model, which refines the selection process by choosing the top queries with the highest semantic similarity from the pool of candidates. In this work, we focus on exploring the robustness issue in the application of QR methodologies, and more specifically, refining the design and implementation of this pipeline to benefit smaller or emerging businesses in regions characterized by multilingual or complex language contexts.
Query reformulation with LLM A more recent idea of QR is to acquire related queries using broad, open-domain knowledge rather than relying solely on in-domain customer data (Jagerman et al., 2023; Wang et al., 2023a; Chen et al., 2023), which has become feasible due to the rise of large language models (LLMs). The implementation falls into two categories: A) use LLMs in the inference stage following a traditional generative paradigm for QR (Zhou et al., 2023; Hao et al., 2022; He et al., 2016; Zuo et al., 2023), e.g., query expansion and query rewrite. This approach has been adopted in E-commerce search to improve recall of long-tail queries (Peng et al., 2023). B) use LLMs for training data augmentation (Wang et al., 2023b), e.g., generate hard negatives. However, the first approach does not fit the purpose of behavioral signal enrichment as it is hard to force rewritten query to have rich behavioral signal. And we will explore how LLM-based data augmentation could dynamically integrate with the hard negative mining used in this paper in future work.
3. Problem formulation
In this section, we present a comprehensive formulation of the query reformulation problem within E-commerce search tasks. We begin with a set of search queries, denoted as $Q$, collected over a specific time frame. We divide this set into distinct subsets: $Q=Q_{r}+Q_{i}$, where $Q_{r}$ denotes behavior-rich queries, $Q_{i}$ denotes behavior-impoverished queries.
For the Query Reformulation (QR) problem, our objective is to establish mappings for each behavior-impoverished query, $q_{i}\in Q_{i}$, such that $q_{i}:\mapsto\{q_{i1},...,q_{iN}\}$, where $q_{ik}\in Q_{r}$, and each query mapping $(q_{i},q_{ik})$ is considered to be semantically relevant. Given the absence of a universal gold standard for defining the semantic relevance of query pairs, we also consider a sub-problem to derive a real-world relevance estimation on the query pair space $Q_{i}\times Q_{r}\mapsto s$, where $s$ denotes a relevance score.
4. Method
To ensure both computational efficiency and accuracy of QR inference, we follow the paradigm introduced by previous research (Huang et al., 2023) and decompose the computation into two stages. In the initial retrieval stage, a bi-encoder (Reimers and Gurevych, 2019) model is employed to generate embeddings for each query in both $Q_{r}$ and $Q_{i}$, constructing a small candidate pool through K-nearest neighbor (KNN) search. In the following re-ranking stage, a cross-encoder (Devlin et al., 2018) model estimates relevance score $s$ for all pairs of a query and its corresponding candidates, based on which only relevant candidates are retained.
In this section, we elaborate on the enhancements introduced to the training of models mentioned above. Our refinements encompass the data mining process to generate training dataset, the model training objective, and an integration of Approximate Nearest Neighbor Negative Contrastive Estimation (ANCE) (Sohn, 2016) into both the retrieval and re-ranking model training processes.
An illustrative representation of the QR pipeline in both model training stage and in online inference system is shown in Figure 1. We highlight the refinements to model training proposed in this paper with bordered text boxes.
Figure 1. An illustration of a) QR pipeline in model training stage (hollow arrows), b) QR pipeline in online inference system (solid arrows), c) data artifact used in training and inference (text box without borderline), d) our refinements to the QR pipeline (text box with borderline).
4.1. Mining query-query training data
The training of bi-encoder and cross-encoder models relies on positive and negative query-query pairs derived from mining customer behavior data (i.e., search log data). This subsection introduce a data mining strategy designed to mitigate false positive cases. The initial step involves extracting the behavior data (i.e., products and their corresponding purchase counts) of each query $q_{i}$, denoted as:
(1) $B(q_{i}):\{p_{i1}:n_{i1},\ p_{i2}:n_{i2},...\}$
where $p_{ik}$ represent a product which has been purchased under query $q_{i}$ and n signifies the corresponding number of purchase. This dataset exclusively consists of queries from $Q_{r}$ as other queries lack a sufficiently strong signal to form positive and negative labels. The behavior data is further refined as $B_{filtered}(q_{i})=\{p_{ik}|n_{ik}\geq P\}$ to filter out the query-product pairs with less than P purchase to reduce noise in the dataset. Finally, the pair $(q_{i},q_{j})$ is retained if $|B_{filtered}(q_{i})\cap B_{filtered}(q_{j})|\geq C$ where ||$\cdot$|| ⋅ | represents the cardinality of a set.
The purpose of the QR mechanism is to facilitate the sharing of behavioral signals between queries. However, it is crucial to mitigate the inadvertent introduction of irrelevant query-product associations into downstream search applications. The key principle is to ensure that, for any query $q_{i}$, only $q_{ik}$ relevant to the shopping intent of $q_{i}$ is retained.
One approach to enforce this constraint in the models’ practice is to prevent the occurrence of ”false positives” in the model training data. For instance, the queries “furniture” and “dining table” share numerous associated products, and are usually considered related in many behavioral mining algorithms. However, borrowing behavior of the query “furniture” to the query “dining table” would inevitably relate products like “TV cabinet” to “dining table”. A prior study (Huang et al., 2023) introduced an advanced positive labeling approach that quantifies the similarity between $q_{1}$ and $q_{2}$ as
(2) $\dfrac{|PP(q_{1})\cap PP(q_{2})|}{\mathrm{Min}(|PP(q_{1})|,|PP(q_{2})|)}\cdot\dfrac{|PP(q_{1})\cap PP(q_{2})|}{|PP(q_{1})\cup PP(q_{2})|}$
with $PP(q)$ representing the set of products purchased under query $q$. Only the top X(i.e., a pre-defined hyper-parameter) pairs associated to each query according to the similarity are defined as positive.
In this paper, we refine the similarity measurement as follows.
  1. (1)
    Normalize the number of purchase in $B_{filtered}(q_{i})$ as $\hat{n_{ik}}=n_{ik}/\sum_{m=1}^{x}n_{im}$.
  2. (2)
    Calculate Jensen–Shannon divergence (JSD) between normalized purchased product count distribution, denoted as $JSD(dist(q_{1})||dist(q_{2})$, where $dist(q_{1})=\{\hat{n_{11}},\hat{n_{12}},...,\hat{n_{1x}}\}$, $dist(q_{2})=\{\hat{n_{21}},\hat{n_{22}},...,\hat{n_{2y}}\}$.
The improved measurement alleviates a common issue where a pair might receive a high similarity score despite having significantly different distributions in their product purchase counts (as shown in Figure 1(a)). Under this refined measurement, only pairs that share not only the same set of purchased products but also a similar distribution can achieve a high similarity score (as shown in Figure 1(b)). To illustrate the impact of this enhancement, we provide examples of the top-1 similar query pairs $(q_{i},q_{i1})$ under both the previous and the new measurement methods in Table 1. It becomes evident that the new measurement yields more suitable query pairs and eliminates undesirable cases, such as ”google pixel watch” and ”pixel watch case.”
(a) Queries with different semantic meaning
(b) Queries sharing identical semantic meaning
Figure 2. The distribution of number of purchase on product under different queries.
Table 1. Example of top-1 similar query pairs derived with the previous similarity metric (Equation 2) and the proposed metric based on JS-divergence.
$q_{i}$
Translation
$q_{i1}$ (previous)
Translation
$q_{i1}$ (proposed)
Translation
google pixel watch
N/A
pixel watch ケース
pixel watch case
pixel watch
N/A
usb bluetooth 5.0
N/A
bluetooth tplink
N/A
tp-link bluetooth usb bluetooth 5.0
N/A
pdp コントローラー
pdp controller
pdp
N/A
pdp コントローラー switch
pdp controller switch
n極対応 延長コード
n-pole compatible extension cord
n極対応
Compatible with n-pole
延長コード n極対応
n-pole compatible extension cord
To further mitigate the occurrence of false positives, we take into account the asymmetrical nature of $(q_{i},q_{ik})$ pairs in the training data. It is noteworthy that $(q_{i},q_{ik})$ forms a suitable pair in the context of QR does not necessarily imply that $(q_{ik},q_{i})$ is an appropriate pair. In the KNN search-based retrieval task, it is infeasible to finetune the model to differentiate between the positions of queries within a pair, as the embedding calculation only takes one query at a time as input. However, in the re-ranking task, the cross-encoder model can be trained in a position-aware mode. Therefore, we use separate similarity measures for retrieval and re-ranking tasks to fully exploit the inductive bias of each model and ultimately address relevance defects. First, JSD is used for retrieval as a symmetric optimization objective would stabilize the bi-encoder training process. Second, consider the calculation
$JSD(dist(q_{i})||dist(q_{ik}))\propto\\ KLD(dist(q_{ik})||dist(m_{(i,ik)}))+KLD(dist(q_{i})||dist(m_{(i,ik)})))$
where $dist(m_{(i,ik)})=(dist(q_{ik})+dist(q_{i}))/2$. The first term emphasizes the distribution divergence with respect to the space defined by products associated to $q_{ik}$, which is more sensitive on false-positive cases where major products that compose $dist(q_{ik})$ is out of the distribution represented by $dist(q_{i})$; while the second term focuses on the other way around, potentially leading a departure from identifying false positives. Therefore, we only retain the first term of JSD in similarity measurement for the re-ranking task so that the result is contingent solely on whether the products associated with $q_{ik}$ are related to $q_{i}$.
4.2. Query normalization and training data augmentation
In the realm of E-commerce, the volume of behavioral data for queries and products within the same category can fluctuate significantly, with differences of up to tenfold across various countries and languages. Successful QR practice in Amazon product search have traditionally leveraged an extensive dataset, encompassing over one billion query-query pairs derived from approximately thirty million unique queries. However, when dealing with countries with a small traffic and featuring a unique language, the availability of data can be notably limited. To address this challenge, we advocate the use of query normalization-based data augmentation in mining query-query data, considering the scale of available data required for effective model training.
In developing our QR framework for non-English and low-traffic markets, we employ a text-normalization pipeline of preprocessing steps, including stop word removal, stemming, script type normalization (e.g., translation between different spellings of the same word, as illustrated by ”こども” to ”子供” in the Japanese case), tokenization (e.g., converting ”子供マスク不織布” to ”子供_マスク_不織布”), and token sorting (e.g., reordering ”子供_マスク_不織布” to ”マスク_子供_不織布”). It is imperative to note that this process excludes the handling of identified named entities, such as brands or media titles, in order to prevent any inadvertent introduction of defects.
Recall that in the construction of the query-query dataset, our approach involves two filtering steps aimed at reducing noise. These steps exclude query-product pairs with a low number of purchases and query-query pairs with a limited number of co-purchased products. As a consequence, a considerable number of relevant query-query pairs are eliminated due to the sparsity of behavioral signals. The introduction of query normalization offers a significant mitigation to this issue. In this study, it substantially increases the count of query-query pairs from about 9 million to about 348 million. Furthermore, it also expands the number of unique queries with low request traffic (i.e., contributing to the last one-third of traffic) covered by the behavioral dataset, elevating it from about 166 thousand to more than 3 million. Please note that this number is derived from the entire query set rather than $Q_{r}$.
Furthermore, it has come to our attention that the similarity scores derived from behavioral data between each member within a group of queries sharing the same shopping intent but slightly different representations can exhibit significant disparities. We illustrate this phenomenon in Table 2, where we present the similarity scores computed between the aggregated behavior of a group of queries all centered around the same shopping intent (in this case, masks for children made of non-woven fabric) and the individual behavior of each query within the group. The queries in this group only differ in terms of tokenization (i.e., the presence or absence of space) and token order, but the similarity scores range from 0.18 to 0.71 and 0.22 to 0.62, respectively. Hence, we contend that clustering query behavior through normalization is a necessary step to acquire similarity scores that more accurately reflect the semantic distinctions between queries.
Table 2. An example in Japanese of query similarity disparity within group. In both groups, the queries convey the same meaning, referring to ”masks for children made of non-woven fabric”, but are in different representations with respect to script-type and tokenization (e.g., the text in gray boxes all refers to ”for children”). The normalized representation of the queries within each group is identical.
Group 1
Query Score
子供_マスク_不織布 0.18
マスク_不織布_子供 0.71
マスク_子ども_不織布 0.44
マスク_不織布_こども 0.30
Group 2
Query Score
子供用マスク_不織布 0.62
マスクこども用不織布 0.29
ことも用マスク不織布 0.22
マスク_子供用_不織布 0.40
4.3. Model training with ANCE mining
Previous research has introduced the concept of Approximate Nearest Neighbor Negative Contrastive Estimation (ANCE) (Sohn, 2016) to address the inefficacy of using in-batch negative in contrastive learning. ANCE entails the incorporation of hard negatives, which are selected from the entire dataset for each query within a batch. This selection process is carried out through a sequence of retrieval model inferences and KNN indexing. In the context of QR, we define the set of hard negatives, denoted as $H(q_{i})$, as the queries that are retrieved for $q_{i}$ but do not share any co-purchases with $q_{i}$.
We opt to utilize $B(q_{i})$ instead of $B_{filtered}(q_{i})$ when constructing the co-purchase dataset for ANCE. This choice is made to minimize the occurrence of strongly relevant pairs with sparse signal being inadvertently misused as hard negatives. We put forth distinct strategies to harness ANCE in the training of retrieval and re-ranking models. For the retrieval model, we employ a self-learning strategy, wherein the bi-encoder is iteratively finetuned with the ANCE dataset it generates. Conversely, for the re-ranking model, we adopt a learning-from-teacher strategy. Under this approach, the cross-encoder is finetuned using the ANCE dataset generated by the bi-encoder model. The ANCE mining and model continuously finetune process is described in Figure 3.
Figure 3. The ANCE mining and model continuous finetune process. Specifically, the filtered dataset is used for initial training of retrieval (bi-encoder) and reranking (cross-encoder) models. Then the bi-encoder model is iteratively finetuned using ANCE data derived by self-retrieved pairs subtracting pairs with co-purchase. At last, the cross-encoder model is finetuned with the ANCE data from the bi-encoder.
4.4. Refined training objectives
We use $s_{BE}(q_{i},q_{j})$ to represent the cosine similarity between L2 normalized embedding of $q_{i}$ and $q_{j}$, and $s_{CE}(q_{i},q_{j})$ to represent the similarity score given by a cross-encoder model. For retrieval task, the general training objective under the contrastive learning with in-batch negative framework is defined as
(3) $\mathcal{L}_{\mathrm{retrieval}}=-1/N\cdot\Sigma_{i,k}\frac{\exp{s_{BE}(q_{i},q_{ik})/\tau}}{\Sigma_{j\in\mathrm{batch}}\exp{s_{BE}(q_{i},q_{j})/\tau}}$
where $N$ is the number of positive query pairs, $\tau$ is a temperature parameter.
The combination of the contrastive learning framework and the infoNCE (Oord et al., 2018) (NT-XENT (Sohn, 2016)) loss has demonstrated superior effectiveness compared to point-wise training using loss such as mean-squared error (MSE) and cross-entropy (CE). However, this objective still exhibit a limitation in contrast to MSE or CE, as it treats every query pair uniformly and does not leverage the similarity score. While this characteristic does not hinder its performance in classic unsupervised contrastive learning tasks, where the positives generated through data augmentation are robust, it poses challenges in the context of QR, where the positives derived from data mining are prone to being false positives. Assigning equal importance to each pair can lead to suboptimal results. In this paper, we refine the loss formulation by adding a sample importance term as
(4) $\mathcal{L}_{\mathrm{retrieval,hard}}=-1/N\cdot\Sigma_{i,k}I_{JS}(q_{i},q_{ik})\cdot\\ \frac{\exp{s_{BE}(q_{i},q_{ik})/\tau}}{\Sigma_{j\in\mathrm{batch}}\exp{s_{BE}(q_{i},q_{j})/\tau}+\Sigma_{j\in\mathrm{sample}(H(q_{i}))}\exp{s_{BE}(q_{i},q_{j})/\tau}}$
where $I_{JS}(q_{i},q_{ik})$ equals to $1-JSD(dist(q_{i})||dist(q_{ik}))$. By doing so, the robustness of the training process is bolstered across varying levels of available resources. On one hand, in the presence of abundant training data, the impact of less pertinent pairs is mitigated, allowing for more focused learning. On the other hand, in scenarios with limited supervised resources, pairs that may not be strictly relevant-yet still bear some relevance-can still make a valuable contribution to the model’s training.
For re-ranking task, we use the cross-entropy loss
(5) $\mathcal{L}_{\mathrm{re-ranking}}=-I_{KL}(q_{i},q_{j})\cdot\log s_{CE}(q_{i},q_{j})-\\ (1-I_{KL}(q_{i},q_{j}))\cdot\log(1-s_{CE}(q_{i},q_{j}))$
where $I_{KL}(q_{i},q_{j})$ equals to $-KLD(dist(q_{j})||(dist(q_{i})+dist(q_{j}))/2$.
For cross-encoder training with hard negative, we group the pairs associated to each $q_{i}$ to form a positive set and employ pairwise loss rather than pointwise loss. The necessity for training in a pairwise manner arises particularly in the case of hard negatives, where positive and negative samples often exhibit only nuanced differences in their text representations. Pairwise training can enable the model to capture the supervised signal with greater precision, reduce variance during training, and thus, enhance the overall training efficacy. Specifically, we use circle loss (Sun et al., 2020), which is designed to direct the model’s attention toward ensuring that the logit for positive pairs surpasses that of negative pairs, rather than exclusively concentrating on the similarity score.
(6) $\mathcal{L}_{\mathrm{re-ranking,hard}}=\log(1+\Sigma_{j\in H(q_{i})}\exp s_{CE}(q_{i},q_{j}))\cdot\\ (1+\Sigma_{q_{ik}}\exp(-s_{CE}(q_{i},q_{ik})))$
5. Experiment
In this section, we present our experimental setup and results. Our experiments were conducted in the Amazon product search, with the primary goal of assessing the scalability and extensibility of our query reformulation techniques. In addition to offline evaluations, we tested our proposed QR solution in two downstream online applications to improve behavior-based search ranking features and optimize search Ads matching through A/B testing.
5.1. Dataset generation
We construct training and testing dataset with one-year anonymized Amazon search data. We first extract query-product pairs from anonymized historical customer behavior data. Then we perform self-join of these query-product pairs on product, and aggregate the joined result by product to obtain the query co-purchase data $D$. At last, we select 5,049 unique queries from $D$ and extract all query pairs containing the selected queries as testing data. The remaining pairs are divided into training and validation data, using a 90%/10% ratio. The statistics of the generated dataset is shown in Appendix A.
5.2. Backbone models
In order to boost training efficacy and efficiency, we developed both bi-encoder and cross-encoder models upon pre-trained language models (PLMs). These models employed a BPE tokenizer with a 300K vocabulary, trained using Amazon’s internal product data via the sentencepiece package (Kudo and Richardson, 2018).
5.3. Offline evaluation criteria
We employ distinct evaluation metrics for the retrieval and re-ranking tasks. For retrieval, we calculate the recall@100 and recall@1000 for both the top-3 paired queries and all queries paired to the given query in the testing data. This separation into top-3 and all queries is motivated by their significance in reflecting the model’s performance in downstream ranking and matching applications, respectively; while evaluating both recall@100 and recall@1000 aims to provide intuitions for determining the number of retrieved candidates. The paired queries are ranked based on the behavior-based similarity score, as described in Section 4.1. For queries with more than 100 or 1000 paired queries, we have slightly adjusted the recall formulation as follows
(7) $\frac{|q_{knn}\cap q_{paired}|}{|q_{paired}|}\rightarrow\frac{|q_{knn}\cap q_{paired}|}{min(|q_{paired}|,|q_{knn}|)}$
where $q_{knn}$ and $q_{paired}$ denote the set of retrieved queries and paired queries of q. Besides, we calculate both micro-recall, which represents the average recall of each query in the testing dataset, and macro-recall, which is derived by dividing the total number of retrieved paired queries by the total number of all paired queries.
For the re-ranking aspect, we evaluate each query using NDCG@3 (Normalized Discounted Cumulative Gain) with the behavior-based similarity score of all paired queries in the testing data. In this evaluation, we also report NDCG@3 for the hard negative scenario, which involves the union of all retrieved queries and paired queries in the calculation. In the hard negative scenario, the score for unpaired queries is set to zero. This evaluation offers a more precise measure of the re-ranking model’s performance in downstream applications, as it closely simulates the context of re-ranking followed by retrieval.
5.4. Offline experiment result
We train a series of bi-encoder model and cross-encoder model, separately for retrieval and re-ranking tasks. The proposed approaches are progressively added upon a baseline model to study the individual benefit brought by each approach. The datasets used for training were constructed through query normalization-based augmentation. Additionally, we tested the models implemented under the current QR solution in Amazon, referred as ”baseline_previous”, whose only difference from baseline was that it employed the criteria defined by Equation 2 for mining positive labels.
Retrieval - The experimental results are detailed in Table 3. The baseline model (index rt2) was trained using contrastive loss (Equation 3). For simplicity, we will analyze the findings specifically for recall@100 in relation to the top-3 queries (in bold font), as the results for all metrics follow a similar trend. Notably, the inclusion of the sample importance term elevated recall@100 by 5.4% (index rt3 vs. rt2). This improvement was further augmented by the addition of ANCE, resulting in an incremental gain of 7.9% (index rt4 vs. rt3). Additionally, through successive iterations of the ANCE dataset and continuous model fine-tuning, recall@100 increased to 0.777 in round 2 (index rt5) and 0.798 in round 3 (index rt6). Note that we only perform three rounds of ANCE as we observed marginal benefit with more rounds. The outcomes suggest that the combined utilization of the sample importance term in the loss function and ANCE substantially enhances the retrieval model’s performance.
Table 3. Offline evaluation of retrieval models.
Index Model Recall (top-3) Recall
@100 @1000 @100 @1000
micro macro micro macro micro macro micro macro
rt1 Baseline_previous 0.40798 0.41115 0.58694 0.59376 0.2747 0.20151 0.44657 0.34673
rt2 Baseline 0.59029 0.59322 0.92091 0.9304 0.47014 0.34508 0.87407 0.74717
rt3 +sample importance 0.64414 0.65069 0.92369 0.93217 0.51185 0.37882 0.86913 0.74003
rt4 +ANCE (round 1) 0.72323 0.72559 0.91881 0.9231 0.6075 0.46382 0.87634 0.74752
rt5 +ANCE (round 2) 0.77664 0.77706 0.93883 0.94461 0.65705 0.48759 0.89501 0.76242
rt6 +ANCE (round 3) 0.79837 0.79949 0.94361 0.94968 0.68076 0.51591 0.9038 0.77851
Re-ranking - the model performance is presented in Table 4. The baseline model was trained using cross-entropy loss. Surprisingly, our findings revealed a rather unexpected outcome: as the model’s focus shifted towards hard negative samples through the incorporation of ANCE and the adoption of circle loss, the model’s performance exhibited a gradual decline in the general scenario, while showing an improvement in the hard negative scenario. This led us to hypothesize that an emphasis on hard negatives during training might result in overfitting. Moreover, this finding prompted a reconsideration of how re-ranking models should be evaluated. Relying solely on paired queries for assessment may not suffice, as it risks producing misleading comparisons among different models. This is particularly pertinent because the practical use of the re-ranking model in the end-to-end inference process mirror hard negative scenarios more closely. Our stance is further substantiated by the results of human auditing, detailed in Section 5.5. Here, we observed that evaluations in hard negative contexts correlated more closely with the outcomes of relevance auditing, underscoring the practical applicability of our model in realistic settings.
Table 4. Offline evaluation of re-ranking models.
Index Model NDCG@3 NDCG@3
(general) (with hard negative)
rr1 Baseline_previous 0.81487 0.30102
rr2 Baseline 0.89459 0.34844
rr3 +ANCE 0.84462 0.64612
rr4 +pairwise loss 0.79172 0.67733
5.5. Offline query-query relevance auditing
Conducting A/B tests online demands considerably more resources than offline analysis and entails the risk of deploying models of inadequate quality into production. Hence, establishing an interpretable pattern that bridges the gap between offline findings and actual production results is immensely beneficial.
Unfortunately, the quantitative metrics such as NDCG and Recall fail to offer this interpretability as they solely provide empirical insights and lack the depth of an explanatory gold standard. Therefore, this section narrows its focus to query relevance evaluation, emphasizing a human-audited approach over evaluating the model’s compatibility with the testing dataset extracted from historical customer behavior. The auditing protocol designed to assess the semantic relevance between queries within the scope of QR is outlined in Appendix B. In summary, we categorize the query-query relevance into three classes - strict relevant, somewhat relevant, and irrelevant.
We extracted a sample of 400 query pairs from the retrieval dataset and conducted a human audit. The assessment of models used for experiment on the audit data is detailed in Table 5. The model performance was measured through the area under the Receiver Operating Characteristic curve (AUROC) and Spearman correlation.
Table 5. Model evaluation based on human audit.
Model AUROC AUROC Spearman
index (strictly relevant) (not relevant) correlation
Retrieval
rt1 $0.51263$ $0.53949$ $0.0373$
rt6 $0.78956$ $0.7481$ $0.4754$
Re-ranking
rr2 $0.71178$ $0.57035$ $0.0572$
rr4 $0.77072$ $0.71539$ $0.4254$
From the findings presented in Table 5, a noticeable distinction between the rt1 and rt6 models, as well as between rr2 and rr4 models emerges. The rt1/rr2 model showcased an inability to discern between relevant and irrelevant queries in the hard negative scenario, as indicated by both AUROC and Spearman correlation measures, whereas the rt6/rr4 model exhibited an opposing performance. This outcome suggests that the proposed techniques fostered robustness of QR through bolstering the model’s capacity in accurately modeling query relevance.
In addition, we noticed that the retrieval model rt6 outperformed the re-ranking model rr4 in terms of both AUROC and Spearman correlation metrics. We attribute this disparity to the distinct training objectives of the two models (i.e., JSD for retrieval and KLD for re-ranking). Furthermore, our auditing protocol is more in tune with the requirements of retrieval, which might explain the observed differences in performance. Please refer to Section 4.1 for a detailed exposition.
5.6. Online experiment
In this section, we report the A/B test result of QR applications in Amazon product search. We focus on two applications. One is to enhance the product search ranking feature. Specifically, we aim at the customer behavior-based query-product level feature, denoted by $f(p,q)$. With a mapping $q_{i}:\mapsto\{q_{i1},...,q_{iN}\}$ given by QR and $f(p,q)$, we augment the targeted feature by
(8) $\hat{f(q_{i},p)}=\alpha f(q_{i},p)+(1-\alpha)\cdot\beta\cdot\frac{1}{N}\Sigma_{k}f(q_{ik},p),$
where $\alpha$ and $\beta$ are hyper-parameters. In this application, we conducted two-week A/B tests in different markets, characterized by Japanese, Hindi, and English, respectively, to evaluate the QR implementation introduced in this study. Specifically, for the Japanese case, we conducted two experiments successively with the implementation suggested by the previous work (rt1+rr1) and our proposed implementation (rt6+rr4) as the treatment, respectively, and the system without any QR solution as control; for the Hindi case, we tested the proposed implementation (rt6+rr4) as treatment and the system without QR as control; for the English case, we tested the proposed implementation (rt6+rr4) as treatment while the previous work (rt1+rr1) as control. The results are detailed in Table 6, revealing that the proposed implementation resulted in a considerable improvement in both revenue and conversion in all tested markets. By comparing the previous implementation versus the new implementation in the Japanese (i.e., rt6+rr4 row verses rt1+rr1 row) and English (i.e., rt6+rr4 row) cases, we can further assert that the proposed methods on model training data mining (such as positive query pair labeling and similarity scoring), training objectives (specifically the addition of a sample importance term), and training strategies (such as employing ANCE) are pivotal components for the success of QR practice in terms of scalability and extensibility.
The second application pertains to product search Ads matching. In our experiment performed in Japanese market, we included the Ads matched for $\{q_{i1},...,q_{iN}\}$ directly into the match set of $q_{i}$. The experiment encompassed both QR implementations, one with the proposed techniques (indexed as rt6 and rr4), and another without them (indexed as rt2 and rr2) through separate two-week A/B tests. Table 6 presents the results, indicating that the model with hard negative-tailored adjustments exhibited a noteworthy gain in both Ads revenue and click-through-rate (CTR), while the model without these adjustments showed only a marginal gain. This outcome strongly suggests that effectively handling hard negatives is a crucial component for the successful practice of QR.
Table 6. Online A/B test results of different applications on markets characterized by different languages.
Product search ranking
Language Control Treatment Revenue Conversion
Japanese w/o QR rt1+rr1 $-0.05\%$ $-0.16\%$
w/o QR rt6+rr4 $+0.14\%$ $+0.08\%$
Hindi w/o QR rt6+rr4 $+0.29\%$ $-0.01\%$
English rt1+rr1 rt6+rr4 $+0.08\%$ $+0.06\%$
Product search Ads matching
Language Control Treatment Ads Revenue Ads CTR
Japanese w/o QR rt2+rr2 $+0.06\%$ $+0.01\%$
w/o QR rt6+rr4 $+0.36\%$ $+0.36\%$
Additionally, we would like to clarify that the experiments conducted for both applications only introduced changes for queries with low traffic, specifically for those representing the last one-third of overall traffic in Amazon’s product search. The reported numbers were derived from the entire traffic analyzed by Amazon’s in-house A/B test system. As a result, the actual impact on queries with low traffic is anticipated to be more substantial than what is being reported.
6. Conclusion
In this study, our primary focus centers on query reformulation (QR) as a remedy to address the limited behavioral data available for queries with historically low customer engagement within the realm of E-commerce. We have identified inherent issues within the existing QR solution, particularly emphasizing challenges related to its scalability and extensibility when applied to smaller markets that lack adequate training data and possess unique language characteristics.
Our investigation was prompted by the inability to replicate the success of the existing QR implementation, proven effective in large Enligsh-markets, in other smaller or emerging markets. To resolve these issues, we have introduced various approaches aimed at enhancing the training of models used in the QR pipeline. These enhancements were focused on refining behavior data mining, as well as optimizing the training strategies and objectives of the models. Through comprehensive A/B testing on two online applications in Amazon product search - specifically, product search ranking and product search Ads matching - we confirmed the necessity and efficacy of these approaches.
Appendix A Statistics of dataset generated in this study
Table 7. Dataset generated in this study.
#unique queries
#positive pairs
#negative pairs
Training+ Validation
5,107,314
339,895,694
N/A
Testing
5,049
371,238
N/A
ANCE (round 1)
5,107,314
339,895,694
743,736,086
ANCE (round 2)
5,107,314
339,895,694
701,336,588
ANCE (round 3)
5,107,314
339,895,694
684,852,381
The statistics of the generated dataset is shown in Table 7. It is worth mentioning that $(q_{i},q_{j})$ and $(q_{j},q_{i})$ are counted as two separate pairs in the counting. We also provide the information on the ANCE dataset. Notabally, the ANCE dataset demonstrates a trend toward convergence across iterations. In the second round, it retains 33.8% of hard negative query pairs from the first round, while in the third round, it retains 64.6% of the pairs from the second round. We only perform three rounds of ANCE as the empirical result shows marginal benefit with more rounds.
Appendix B Human audit protocol
The human auditor is tasked with categorizing each query-query pair into one of three following classes.
Strict relevant - Query 1 and query 2 are different representations of the same shopping intent (i.e., customer would expect the search result of query 1 and query 2 to largely overlap with each other).
Somewhat relevant - The shopping intent of query 1 and query 2 have no conflict but are not exactly the same, for example, query 1’s shopping intent is much broader and entirely cover query 2’s, or the opposite case (i.e., customer expect to find item related to query 2 when search for query 1, or vice versa. But the overlap between search result of query 1 and 2 is only a small portion of the search result of query 1 or 2).
Not relevant - Query 1 and query 2 have some related components, but share conflict shopping intent in either of 1) Product Type, 2) Model, 3) Age/gender group, 4) Size/color.